A train traveling at 100 kmph overtakes a motorbike traveling at 64 kmph in 40 seconds. What is the length of the train in meters?
Answer by khwang(438)

Let the length of the train be L meters.
Since the time for the train to overtake the motorbike is
L/(100-64) hrs (why?)
Now we have L/(100-64) = 40/3600. (1 hr =3600 sec)
So, L= 40/100 = 0.4 meters
Kenny